OpenStudy (anonymous):

write equation in slope intercept form walk through (4,4) slope =2

OpenStudy (anonymous):

is is y=mx+b?

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

yes

OpenStudy (goformit100):



OpenStudy (anonymous):

y=2x+b, plug in (4,4) to find b 4=8+b, b=-4 so y=2x-4

OpenStudy (anonymous):

how did it turn into 4=8+b?

OpenStudy (anonymous):

they gave you a point (4,4). so y is 4 when x is 4. y=2x+c; 4=2(4)+b implies 4=8+c

OpenStudy (anonymous):

sorry b, not c

OpenStudy (anonymous):

oh i see now thank you.

OpenStudy (anonymous):

ur welcome